A steam era no 09 tle steam at- i a? °C Ass aeud irn ne msk
Answer #1

(a) The efficiency of the engine will be

\eta=1-\frac{T_c}{T_h}=1-\frac{127+273}{427+273}=1-\frac{400}{700}=\frac{300}{700}=0.4286

(b) Output work will be

W=\eta Q=0.4286\times2000=857.143$ J$

(c) Energy exhausted will be

Q'=Q-W=Q-\eta Q=2000-857.143=1142.857$ J$

(d) Let the temperature of hot steam be T_h' then

2\eta=1-\frac{T_c}{T_h'}\Rightarrow 0.4286\times2=1-\frac{400}{T_h'}\Rightarrow \frac{400}{T_h'}=0.143\\ T_h'=\frac{400}{0.143} =2797.2^o$ K$=2797.2-273=2524.2^0$ C$